OpenStudy (anonymous):

URGENTTT!! Find the force a 68.5 kg person would experience while standing on the surface of Mars. Mars has a mass of 6.43 x 1023 kg and a radius of 3.43 x 106m from its center to the surface.

OpenStudy (anonymous):

can someone at least tell me what formula to use?

OpenStudy (anonymous):

You use the gravitational formula \(Fg=mg\). But you don't have the g constant (it will be different since it is on Mars\). To find the constant g, you should use \(g=\frac{GM}{r^2}\), where G is the universal gravitational constant \((6.673 \times 10^{-11})\), M is the mass of MARS and r is the radius of mars.

OpenStudy (anonymous):

have you tried what i told you?

OpenStudy (anonymous):

you have to find \(\LARGE g\) first. using the formula \(\LARGE g=\frac{GM}{r^2}\) where \(\LARGE {G=6.673 \times 10^{-11}, M=6.43 \times 10^{23} kg\\ and\ r=3.43 \times 10^6m}\)

OpenStudy (anonymous):

by the way, \(\Large \color{blue}{Welcome}\ \color{orange}{to}\ \color{purple}{OPENSTUDY! ☺}\)

OpenStudy (anonymous):

thanks! and i still cant find the answer,These are the answers : a). 790 N b). 259 N c). 246 N d). 8.57 x 108 N But no matter what I try my answer is totally different HELP!!

OpenStudy (anonymous):

can you show me your solution?

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

(6.673x10^-11)*(6.43x10^23)/(3.43x10^6)x(2) =2.50188863x10^19

OpenStudy (anonymous):

For the bottom part, it is r squared not r times two

OpenStudy (anonymous):

well yes but it still gives me a long decimal similar

OpenStudy (anonymous):

no it won't, try plugging the numbers carefully

OpenStudy (anonymous):

i dont know what i m doing wrong

OpenStudy (anonymous):

let's try to do it step by step. what's \(6.673x10^{-11} \times 6.43x10^{23}\) ? what's \((3.43 \times 10^{6})^2\)?

OpenStudy (anonymous):

it's close to 'b' and 'c'(i got approximately 250 N)

OpenStudy (anonymous):

i think it c